Data concerning a recent period’s activity in the Prep Department, the first processing department in a company that uses process costing, appear below:

Materials Conversion
Equivalent units in ending work in process inventory 2,180 810
Cost per equivalent unit $ 14.26 $ 6.43

A total of 20,900 units were completed and transferred to the next processing department during the period.

Required:

1. Compute the cost of ending work in process inventory for materials, conversion, and in total.

2. Compute the cost of the units completed and transferred out for materials, conversion, and in total.

(Round your final answers to the nearest whole dollar amount.)
Answer #1

Cost of ending WIP

Material Conversion Total
Cost of ending WIP 2180*14.26 = 31087 810*6.43 = 5208 36295

Cost of Unit transferred out

Material Conversion Total
Cost of transferred out 20900*14.26 = 298034 20900*6.43 = 134387 432421